OpenStudy (anonymous):

The Mean Value Theorem can be applied to f(x)=x^2 on [-2,1]. Find all values of c in (-2,1) satisfying the conclusion of the Mean Value Theorem

OpenStudy (amistre64):

whats the slope between the endpoints?

OpenStudy (anonymous):

I'm not sure, I know have to derive it to f'(x)=2x and then have it f'(c)=2(c) but I don't know what to do from there

OpenStudy (amistre64):

youve got the hardest part figured out, we simply need to compare it to the slope between the endpoints of your interval

OpenStudy (amistre64):

how do we find the slope between 2 points?

OpenStudy (anonymous):

f(b)-f(a)/b-a?

OpenStudy (amistre64):

exactly

OpenStudy (anonymous):

what are the steps though? She didn't explain it well on the notes

OpenStudy (amistre64):

an interval is from a to b: (a,b) so we know what a and b are the function f is stated as x^2 so, what is f(b)-f(a) and what is b-a?

OpenStudy (anonymous):

f(-2)-f(1)/-2-1?.. and I got -3/-3

OpenStudy (amistre64):

not sure if there is a stepping process, its a matter if finding out certain things and putting it all together determine the slope between the endpoints. determine the derivative of f (find f') equate them to determine a value for c such that: f'(c) = (f(b)-f(a))/(b-a)

OpenStudy (amistre64):

f=x^2 f' = 2x [-2,1] -2^2 = 4 1^2=1 4-1 = 3 1--2 = 3 3/3 = 1 is the slope between endpoints when does 2c = 1?

OpenStudy (amistre64):

might have thought to quickly on slope ... b=1, a=-2 1 - 4 = -3 slope is -1

OpenStudy (amistre64):



OpenStudy (anonymous):

Ooh. So then the slope would be equal to the f'(c)?

OpenStudy (anonymous):

And that's how we get -1/2?

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

2c = -1, c = -1/2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (amistre64):

youre welcome :)